you will receive $100 from a zero-coupon savings bond in 2 years. The nominal interest rate is 8.10%.

If the inflation rate over the next few years is expected to be 3.10%, what will the real value of the $100 payoff be in terms of today’s dollars? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

-answer is $94.08..

but my Q is this....

d. Show that the real payoff from the bond [from part (b)] discounted at the real interest rate [from part (c)] gives the same present value for the bond as you found in part (a). (Do not round intermediate calculations. Round your answer to 2 decimal places.)

*present value :

what is present value when real value was $94.08

plz solve this Q
Answer #1

real rate=(1+nominal rate)/(1+inflation rate)-1=1.081/1.031-1=4.850%

Present value=Real value/(1+real rate)^2
=94.08/(1+4.850%)^2
=$85.58